OpenStudy (abbycross167):

37. Which equation describes the proportion: The number of chickens (C) is four times as many the number of horses (H) on the farm.? a. C = 4 H b. C = ¼ H c. H = C + 4 d. C = H + 4 My Answer ~ a 38. Given 3x – y = 8. Solve for x. a. x = (8 - y)/3 b. x = (8 + y)/3 c. x = -8 – 3y d. x = -8 + 3y My Answer~ a

OpenStudy (yese30):

try 37. C

OpenStudy (abbycross167):

what do you mean?

OpenStudy (phi):

you have 4 times as many chickens as horses. for example if you have 1 horse you have 4 chickens you would write the proportion \[ \frac{c}{h}= \frac{4}{1} \\ \frac{c}{h}=4\] For Question 38 3x – y = 8 to solve for x, start by adding +y to both sides 3x -y+y = 8+y on the left side -y+y is zero and 3x+0 is 3x so you have 3x = 8+y now divide both sides by 3 \[ x= \frac{8+y}{3} \]

OpenStudy (abbycross167):

so #37 would be a or b? and #38 would be b?

OpenStudy (radar):

No, there is a difference between A and B.

OpenStudy (abbycross167):

hmm ok, I don't know which one it would be...

OpenStudy (radar):

Look at phi posty.